OpenStudy (parissfrance09):

What is the solution to the equation? please help!! 3q + 5 = –2q + 20 A. q = 3 B. q = 5 C. q = 15 D. q = 25

OpenStudy (anonymous):

Plug in the numbers and see which one works. (In this case, it's 15.) \[3*15+5 = -2*15 + 20\]50 = 50

OpenStudy (parissfrance09):

thank you